Two capacitors with 2.0 F and 3.0 F capacitance, respectively, are connected in series and subjected to a total potential difference of 100 V. Find the (a) total capacitance,


(b) charge stored in each capacitor, and (c) potential difference across each capacitor

1
Expert's answer
2022-03-26T15:31:55-0400

(a) total capacitance



"C=\frac{C_1C_2}{C_1+C_2}""C=\frac{2.0*3.0}{2.0+3.0}=1.2\:\rm F"

(b) charge stored in each capacitor



"q=CV\\
=1.2*100=120\:\rm C"

(c) potential difference across each capacitor



"V_1=q/C_1=120/2.0=60\:\rm V""V_2=q/C_2=120/3.0=40\:\rm V"